OpenStudy (anonymous):

A computer program changes one number into another. The program changes 4 into -5, -2 into -11, and 33 into 24. What will 9 be changed into?

OpenStudy (anonymous):

9 will be changed into 0. The program subtracts 9 from each number (4-9=-5, -2-9=-11, and 33-9=24).

OpenStudy (anonymous):

9 will change into 0